OpenStudy (anonymous):

A CERTAIN SPECIES OF CRICKETS HAS A LINEAR RELATIONSHIP BETWEEN THE AIR TEMPERATURE "T" AND THE NUMBER OF CHIRPS PER MINUTE BY THE CRICKET. tHIS LINEAR REALTIONSHIP IS GIVEN BY ch(T)=4T-160. WHEN THE TEMPERATURE "T" IS 55 DEGREES, FIND THE NUMBER OF CRICKET CHIRPS PER MINUTE. pLEASE COULD SOME ONE WALK ME THROUGH STEP BY STEP HOW TO SOLVE THIS?

OpenStudy (anonymous):

ch(T) gives cricket chirps in terms of T if T = 55 & ch(T) = 4T -160, we just plug in 55 for T (substitution) Therefore: ch(T) = 4*55 - 160 = 170 - 160 = 10 chirps/minute

OpenStudy (anonymous):

THAT DOESNT LOOK RIGHT TO ME, IF YOU 4*55-160, MY ANSWER COMES UP WITH 220-160 WHICH THEN IS 60 FOR FINAL ANSWER RIGHT?

OpenStudy (anonymous):

You are right. Multiplication error :D

OpenStudy (anonymous):

oh ok thanks i dk you answered already